OpenStudy (anonymous):

Help Please! Reduce rational expression, x to 2 power-9 5x-x to the 2 power x-4 _______________muli __________________ div ________ x to 2 power-5x x to 2 power-x-12 x to 2 power -8x+16 Thank you.

OpenStudy (anonymous):

i got 1 but i know that can't be right.

OpenStudy (saifoo.khan):

\[\frac{(x+3)(x-3)}{x(x-5)} \times \frac{-x(x-5)}{(x-4) (x+3)} \times \frac{(x-4)(x-4)}{(x-4)}\]

OpenStudy (saifoo.khan):

wow, that was HUGE.

OpenStudy (anonymous):

yes! so it is 1?

OpenStudy (saifoo.khan):

solving..

OpenStudy (saifoo.khan):

isnt it -(x-3) ? im still checking.

OpenStudy (anonymous):

yes it must be x-3, everything else cancels each other out

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

Right, so it's -x+3

OpenStudy (anonymous):

why is it negative?

OpenStudy (saifoo.khan):

check in the second fraction, i took -x common instead of x. Later the x cancels leaving behind -1.

OpenStudy (saifoo.khan):

So, -1( x-3) -x + 3

OpenStudy (anonymous):

yes, i see... got it, thank you.

OpenStudy (anonymous):

little things like that always screw me up.

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

so i have six more problems to do, if your up to it

OpenStudy (saifoo.khan):

Sure.